OpenStudy (anonymous):

Find the break even point for the profit function P(x) = 2.5x - 15,000. Please show your work.

OpenStudy (anonymous):

2.5x-15,000=0 2.5x=15,000 x=15,000/2.5 x=6,000

OpenStudy (anonymous):

the break even values for a profit model are the values for which you earn 0 (zero) in profit so 2.5x-15,000=0 2.5x=15,000 x=15,000/2.5 x=6,000 so break even point for the profit function P(x) = 2.5x - 15,000 is x which is equal to 6000

OpenStudy (anonymous):

Thank you!